Please post steps for each

adbook ORION Downloadable eTextbook ent FULL SCREEN PRINTER VERSION BACK | NEXT Chapter 06, Problem 65 Your answer is partially correct. Try again. Interactive Solution 6.65 offers a model for solving this problem. A car accelerates uniformly from rest to 21.7 m/s in 7.02 the average power required to accelerate the car if (a) the s along a level stretch of road. Ignoring friction, determine welght of the car is 7.60 x 103 N, and (b) the weight of the car is 1.05 x 104 N (a) Number T15444.096 (b) Number 71870.4 SHOW HINT 10/17/2017
Answer #1

a)

Vo = initial speed = 0 m/s

Vf = final speed = 21.7 m/s

m = mass = weight/g = 7600/9.8 = 775.51 kg

t = time = 7.02 s

Average power is given as

P = (0.5) m (Vf2 - Vo2)/t

P = (0.5) (775.51) (21.72 - 02)/7.02

P = 26010 Watt

b)

Vo = initial speed = 0 m/s

Vf = final speed = 21.7 m/s

m = mass = weight/g = 10500/9.8 = 1071.43 kg

t = time = 7.02 s

Average power is given as

P = (0.5) m (Vf2 - Vo2)/t

P = (0.5) (1071.43) (21.72 - 02)/7.02

P = 35934.88 Watt